OpenStudy (anonymous):

A canoe can go 24 km downstream in 3 hours. The return trip takes 4 hours. What is the speed of the current?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

What would be a system of equations for this problem?

OpenStudy (anonymous):

Here's how we go about, You know that distance/speed = Time ?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Ok. We take 2 speeds into account in such questions. 1 - speed of the boat in still water - let it be x km/h 2 - speed of current - let it be y km/h (which is what we have to find.) In upstream, we go against current so speed = (x-y) km/h, in downstream we go with the current so speed = (x+y) km/h. Now the boat goes 24 km upstream in 3 hours. so the speed of the boat upstream = 24/3 = 8 km/hr. Hence x-y = 8; fine till here?

OpenStudy (anonymous):

oops sorry its downstream; so x+y = 8 :PP

OpenStudy (anonymous):

Yeah! thank you so much!

OpenStudy (anonymous):

so upstream it will be x-y = 6. solve x+y = 8; x-y = 6. Y = speed of stream will be found to be 1 km/h,

OpenStudy (anonymous):

oh goodness! thank you! this helped a lot!

OpenStudy (anonymous):

i hope that tallies with your book answer?

OpenStudy (anonymous):

It's for an online class so I won't know until tomorrow afternoon!

OpenStudy (anonymous):

oh lol ok..so good luck for that :)